# Teaching Transformers Causal Reasoning through Axiomatic Training

### v2

@@ 5.4 Results: Generalization to Complex Causal Scenarios
$\mathbf{MultiEval_{SLR}}$ Dataset. We evaluate our models and other baselines on the challenging $\mathbf{MultiEval_{SLR}}$ dataset since it includes example that are different from training dataset in terms of size and complexity of causal graph. Table 1 summarizes the results for this dataset. While GPT-4 performs best, models trained with RoPE position encodings still achieve strong results, surpassing Gemini Pro and Phi-3 in both zero-shot and multi-shot settings for majority of node lengths.
@@
Table 2: Evaluation with branching factor 1.4. Accuracy of axiomatically trained models with LM baselines on the causal graphs with branching factor 1.4. See Sec 5.4 for details.
@@
Further Ablations. We run further ablations to understand the generalization behavior of our models. In particular, we experiment with generalization to unseen node names (Appendix Table A5), generalization to unseen lengths for graphs with linear chains and random flipping of edges (Appendix Table A4), and generalization to graphs with branching factor of 2 (A6).
@@
Figure 3: Generalizing to longer unseen causal sequences ($>$6 nodes) with random flipping on TS2 and OCC (with NoPE) train sets. OCC-trained models struggle due to limited edge-level variability, while TS2 NoPE consistently performs well. Refer table A4 for complete results
@@ 6 Axiomatic Training for D-Separation Rule
Similar to the transitivity axiom, we now train our model on instances on of d-separation rule from multiple causal graphs and different premise, hypothesis pairs.
@@ 6.1 Training Dataset and Setup
We follow a similar strategy to generate the training dataset as we did for the transitivity axioms. The training dataset consists of graphs with lengths from $[3,6]$ with branching factor in range $[0.6,0.8]$ and uses the same premises from the TS2 training set as we did for transitivity axiom (see Sec 5.1). Given a premise, we create the hypotheses as follows: First, select all pairs of nodes in the graphs $(x_{1},x_{2})$, then select the conditioning set $\mathcal{C}$ from the remaining sets of nodes of size up to 5 nodes at a time. The ground truth labels denote whether $x_{1}$ and $x_{2}$ are d-separated given conditioning set $\mathcal{C}$ for the given causal graph in the premise. From this exhaustive set of hypotheses, we randomly subsample 175k examples.
CLEAR D-Separation task (YN)
Models Accuracy
Llama-3-8b-Instruct 60.0
Llama-3-8b-Instruct Finetuned 70.0
GPT-4 63.33
CLEAR D-Separation task (MC)
Models Accuracy
Llama-3-8b-Instruct 33.0
Llama-3-8b-Instruct Finetuned 50.0
GPT-4 36.67
Table 3: Evaluation on CLEAR dataset (Chen et al., 2024b) We finetune the LMs on our d-separation dataset and evaluate on the CLEAR dataset in a zero-shot setting. We observe a significant increase in performance compared to the baseline, which highlights the efficacy of axiomatic training. For discussion refer Sec 7.1.
@@ 6.2 Evaluation
Building on the success of axiomatic training for the transitivity axiom, we extend our approach to d-separation by introducing structurally more complex scenarios, such as branching and longer chains with random flipping. Unlike transitivity, which primarily involves reasoning over linear chains, d-separation is inherently more challenging due to its dependence on various structural patterns, including colliders, chains, and forks.
We evaluate the trained 67M model on two evaluation settings: longer sequences with random flipping (refer Table A7) and branching (refer Table 4), to cover a range of structural variations. Overall, model trained with RoPE emerges as the best performer, with NoPE based model following as close second. While GPT-4 struggles to perform better than random baselines in both settings, our models trained from scratch perform much better than random baseline, and the performance goes down as size increases.
@@
GPT-4 0.53 0.544 0.62 0.52
Finetuned Results
Llama-3-8b-Instruct 0.474 0.490 0.470 0.482
Finetuned Llama 0.796 0.738 0.718 0.670
@@
Table 4: We evaluate the effectiveness of axiomatic training for d-separation under two training paradigms: training a model from scratch and fine-tuning a pretrained Llama model. Our training setup consists of linear sequential causal chains, along with some variations where edges are randomly flipped. However, we assess model performance on significantly more complex causal graphs featuring branching structures and additional nodes, similar to our transitivity analysis. The base Llama Instruct model (prior to fine-tuning) performs on par with random baselines. In contrast, the fine-tuned model demonstrates a substantial improvement, particularly on branched networks. Unlike our transitivity analysis where GPT-4 significantly outperformed all other models, GPT-4 struggles in this setting, performing no better than the random baseline. Our experiments with a decoder-based model trained from scratch show superior performance compared to baselines.
7 Evaluating on complex tasks with axiomatic finetuning
We trained Llama-3.1-8B-Instruct using supervised fine-tuning on the same axiomatic training data, where the model takes causal graph premises and hypotheses (for transitivity or d-separation) as inputs and predicts ’Yes’/’No’ as output labels. Refer I for more details.
@@ 7.1 Evaluation on CLEAR Benchmark
Chen et al. (2024a) introduced CLEAR, a comprehensive benchmark assessing LLMs’ causal reasoning across 20 tasks, including backdoor adjustment, d-separation and others. It features diverse question types beyond Yes/No, allowing for evaluation on more complex tasks. We test our model, fine-tuned on axiomatic d-separation instances, in a zero-shot setting on CLEAR’s Yes/No (YN) and Multi-Choice (MC) questions. Despite training on YN labels, our model outperformed GPT-4 on MC tasks. Fine-tuning on axiomatic instances of d-separation yielded a 20% improvement over the base model, surpassing GPT-4, while model showed improvement for YN task as well. Despite differences in semantic structure and wordings of finetuning and evaluation instances, our model, exhibited a significant performance gain over the base model. The substantial zero-shot gains across diverse tasks, semanmtic structure and question types, outperforming larger models like GPT-4, suggest that the model effectively applies d-separation rather than relying on spurious associations. Refer Table 3 for results.
@@ 7.2 Evaluation on Corr2Cause Dataset
Llama-3-8b-Instruct 0.18 0.14 0.23 0.67
Llama-3-8b-Instruct 0.20 0.17 0.24 0.70
Finetuned Dsep (175k)
Llama-3-8b-Instruct 0.37 0.28 0.57 0.70
finetuned transitivity (300k)
GPT-4 (from paper) 0.29 0.21 0.48 0.64
Table 5: Evaluation on the Corr2Cause Task from (Jin et al., 2024b). We finetune the LMs on our transitivity and d-separation dataset and evaluate the Corr2Cause dataset in a zero-shot setting. We observe a significant increase in performance compared to the baseline, which highlights the efficacy of axiomatic training. For discussion refer Sec 7.2.
Jin et al. (2024c) proposed a more complex dataset to evaluate models on different causal tasks. Each data instance in the benchmark includes correlational relationships described in natural language for graphs with 3 to 6 nodes; the goal is to infer the truth value of a hypothesis. The hypothesis consists of six different kinds of graphical relationships between pairs of variables: Parent, Ancestor, Child, Descendant, Collider, and Confounder. This task is significantly harder than applying a single axiom. First, one needs to infer the causal graph from the correlation statements. This requires knowledge of d-separation statements and Markov condition Appendix D.1. Then, one needs to use the transitivity axiom to identify the direct effect and indirect effect to identify the children, ancestors, colliders, and confounders in the causal graph.
Comparison with Baselines:
Our results highlight the Llama-3-8b-Instruct Base model’s poor performance on this task, while axiomatic fine-tuning enables it to handle the complexity effectively. Fine-tuning on transitivity and d-separation improved performance, despite their simplicity. Notably, transitivity fine-tuning led to the largest gains, even surpassing GPT-4, likely due to its direct relevance in inferring graph relationships. This demonstrates the potential of our minimalist training setup to tackle complex, language-based causal reasoning tasks.
@@ 8 Discussion and Conclusion
In this paper, we provide a general framework, axiomatic training, to add axioms and simple rules of causality as inductive prior in the ML models, which can then further help in downstream causal discovery and causal inference tasks. We demonstrate their usefulness by training/fine-tuning models using our axiomatic training framework on two simple rules – transitivity and d-separation rules. We also discuss various modelling choices and how they effect the learning and generalization of causal axioms. We observe that a transformer model trained from scratch on a large axiomatic dataset can learn to apply axioms effectively. On causal tasks like graph traversal via transitivity and inferring causal relationships from correlation, small 67M transformers generalize well to unseen complex graphs, often outperforming models like GPT-4, Phi-3, and Gemini Pro. We then demonstrate the usefulness of adding these rules as inductive prior to downstream tasks on two downstream datasets – CLEAR and Corr2Cause. We observe that fine-tuning LLMs using axiomatic training help perform better on these dataset in the zero-shot setting, i.e., having never seen examples from these datasets. Fine-tuning on d-separation and transitivity-based axiomatic instances led to performance improvements. The model fine-tuned on transitivity exhibited the highest performance gain on templates like Parent, where distinguishing direct and indirect relationships is crucial. Similarly, fine-tuning on d-separation instances resulted in performance improvements on templates like idenrifying collider, where identifying colliders is key. Since the Child template contains no "No" labels, its reported F1 score is 0.
Generalization to Logical Reasoning. While our axiomatic training approach focuses on causal reasoning, it can be applied to any formal system such as deductive logical reasoning. Recent work (Saparov et al., 2023) highlights LLMs’ deterioration in performance as reasoning depth increases. Given the similarity between our setup and such tasks, it would be interesting to explore if axiomatic training can improve deductive reasoning in LMs.
Implications for Training Language Models. GPT-4 demonstrates impressive generalization on the causal tasks we evaluated. We hypothesize that axiomatic training may explain GPT-4’s ability to reason over causal graphs, as (noisy) demonstrations of the underlying axioms could be present in its web-scale training data. Meanwhile, models like Gemini Pro and Phi-3 struggle with zero-shot reasoning for causal tasks, such as handling completely reversed chains, suggesting room for improvement. Incorporating causal axiom demonstrations as a part of language models’ pretraining (or finetuning data) could help improve the reasoning of these models so that small language models like Phi-3 can achieve GPT4-like accuracy on causal tasks. Incorporating axiomatic inductive biases could aid the language model’s reasoning abilities for the desired task. For instance, (Papadimitriou & Jurafsky, 2023) propose pretraining language models on synthetic formal languages to incorporate inductive biases to improve performance in learning grammatically diverse languages.
@@
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Additional Results: Role of Data Diversity and Positional Encoding
Appendix B Multi-Shot Prompt
B.1 Cause-Effect Inference Task
@@
Appendix C Results and Analysis
@@
Table A5: Results on node name length generalization. TS1 and TS2 denote Training Data setup 1 and 2 from Section 4 LABEL:sec:. OCC is the third data setup comprising of sequential causal chains. SPE: Sinusoidal PE, LPE: Learnable PE, w/o PE: No PE, RoPE: Rotary Position Embedding. Model remains the same across all setups (67 Million parameter based). For longer node names, NoPE performs best on sequential linear setup. Accuracy metric is used.
@@
Appendix D Preliminaries and Notations
@@
Let $\mathcal{M}=(\bm{X},\bm{U},\mathcal{F})$ be a causal model defined over a set of endogenous variables $\bm{X}$, exogenous variables $\bm{U}$ and the causal relationship between then defined by set of structural equations $\mathcal{F}$ (Galles & Pearl, 1997). Let $\mathcal{G}$ be the causal graph associated with the causal model $\mathcal{M}$ where the nodes $\bm{V}$ in $\mathcal{G}$ correspond to the variables in $\mathcal{M}$ and an edge $V_{i}\rightarrow V_{j}$ between any two nodes $V_{i},V_{j}$ denote the causal relationship between them. The causal relationship of node $X_{i}$ is characterized by the functional relationship $f_{i}\in\mathcal{F}$ s.t., $x_{i}=f_{i}(\bm{pa}_{i},\bm{u}_{i})$. Here $\bm{pa}_{i}$ are the parent of the node $X_{i}$ is the corresponding causal graph $\mathcal{G}$ and $\bm{u_{i}}\subseteq\bm{U}$ are set of exogenous variables influencing the exogenous variable $X_{i}$. In our work, we assume that there are no hidden confounders so we have one exogenous variable corresponding to every endogenous variable i.e. $\bm{u}_{i}=u_{i}$. Each exogenous variable has an associated probability distribution which quantifies the uncertainty in the system i.e. $u_{i}\sim\mathbb{P}(u_{i})$. Thus the joint distribution of the exogenous variable is given by $\mathbb{P}(\bm{U})$. Since any endogenous variable is a deterministic function of other endogenous and exogenous variables the probability distribution corresponding to the endogenous variable is the push-forward of the exogenous variable i.e $\mathbb{P}(\bm{X})\triangleq\mathbb{P}^{\#}(\bm{U})$.
D.1 Definitions
Following the formal definitions provided by (Jin et al., 2024b), we explain the following terminologies:

### v3

@@ 5.2 Results: Generalizing to Complex Causal Graphs
$\mathbf{MultiEval_{SLR}}$ Dataset. We evaluate our models and other baselines on the challenging $\mathbf{MultiEval_{SLR}}$ dataset since it includes examples that are different from training dataset in terms of size and complexity of causal graph. Table 1 summarizes the results for this dataset. While GPT-4 performs best, models trained with RoPE position encodings still achieve strong results, surpassing Gemini Pro and Phi-3 in both zero-shot and multi-shot settings for majority of node lengths.
@@
Table 2: Evaluation with branching factor 1.4. Accuracy of axiomatically trained models with LM baselines on the causal graphs with branching factor 1.4.
@@
Further Ablations. We run further ablations to understand the generalization behavior of our models. In particular, we experiment with generalization to unseen node names (Appendix Table A5), generalization to unseen lengths for graphs with linear chains and random flipping of edges (Appendix Table A4), and generalization to graphs with branching factor of 2 (Tab. A6).
@@
Figure 3: Generalizing to longer unseen causal sequences (>>6 nodes) with random flipping on TS2 and OCC (with NoPE) train sets. OCC-trained models struggle due to limited edge-level variability, while TS2 NoPE consistently performs well. Refer table A4 for complete results
@@ 6 Axiomatic Training for D-Separation Rule
Similar to the transitivity axiom, we now train our model on instances of the d-separation rule.
@@ 6.1 Training Dataset and Setup
We follow a similar strategy to generate the training dataset as for the transitivity axiom. The training dataset consists of graphs with nodes from $[3,6]$ with branching factor in range $[0.6,0.8]$ and uses the same premises from the TS2 training set (see Sec 5.1). Given a causal graph premise, we create the hypotheses as follows: First, select all pairs of nodes $(x_{1},x_{2})$ in the graph, then select the conditioning set $\mathcal{C}$ from the remaining sets of nodes of size up to 5 nodes. The ground truth labels denote whether $x_{1}$ and $x_{2}$ are d-separated given the conditioning set $\mathcal{C}$. From this exhaustive set of hypotheses, we randomly subsample 175k instances.
@@ 6.2 Evaluation
Unlike transitivity that primarily involves reasoning over linear chains, d-separation is more challenging due to its dependence on various structural patterns, including colliders, chains, and forks. Given the difficulty of the task, we evaluate two models: 1) 67M model trained from scratch on the d-separation axiomatic data; 2) Llama-3-8B model fine-tuned on the same axiomatic data. We consider two evaluation settings: longer sequences with random flipping (refer Table A7) and branching (refer Table 3).
Overall, the model trained with RoPE achieves the highest performance among all models trained from scratch, with the NoPE-based model ranking a close second across both complex graph structures. GPT-4, in contrast, struggles to outperform random baselines in both settings. For complex linear graphs, our RoPE-trained model significantly exceeds random baseline performance and clearly outperforms both GPT-4 and Llama-3-8B-Instruct. The axiomatically fine-tuned model performs comparably and maintains its effectiveness even as graph length increases, whereas the RoPE-trained model’s performance declines for longer sequences. For branched causal graphs, although the RoPE-trained model initially performs well, its accuracy quickly drops toward the random baseline as complexity grows. In these more challenging scenarios, the Llama-fine-tuned model remains consistently strong and surpasses GPT-4 across all node sizes, thus highlighting the impact of axiomatic training in this setup.
@@
GPT-4 0.53 0.54 0.62 0.52
fine-tuned Results
Llama-3-8B-Instruct 0.47 0.49 0.47 0.48
fine-tuned Llama 0.79 0.74 0.72 0.67
@@
Table 3: Effectiveness of axiomatic training for d-separation under two training paradigms: training a model from scratch and fine-tuning a pretrained Llama model.
7 Axiomatic Fine-Tuning for Complex Tasks
We now focus on causal reasoning benchmarks such as Corr2Cause and CLEAR and show how axiomatic fine-tuning can help improve causal reasoning of pretrained language models. We fine-tune Llama-3-8B-Instruct using supervised fine-tuning on our axiomatic training data (described in Sec. 4.1). Fine-tuning an existing model avoids any restriction on tokens and language style; the axiomatically fine-tuned model can be applied to any general text. We use Llama-3-8B-Instruct as the baseline to assess the improvement through axiomatic fine-tuning. We also include GPT-4 as a baseline, given its demonstrated strength in reasoning tasks.
@@ 7.1 Evaluation on CLEAR Benchmark
Dataset and Evaluation Setup. CLEAR (Chen et al., 2024) provides a comprehensive benchmark assessing language models’ understanding of causal graphs through 20 diverse tasks, including d-separation. It has different question types like Multi-Choice and Yes/No type questions (Refer Fig. 1 for an example instance).
We test our model, fine-tuned on axiomatic d-separation instances, in a zero-shot setting on CLEAR’s Yes/No (YN) and Multi-Choice (MC) questions. Note that there are multiple differences, including in structure and wording, between the fine-tuning data and the test benchmark. In addition, our fine-tuning data does not include any MC questions, only YN questions. Our goal is to check whether a axiomatically fine-tuned Llama model would generalize zero-shot to new kinds of causal questions.
Results. Table 5 shows the results. Our axiomatically fine-tuned Llama model provides a significant improvement in accuracy over the base model, for both YN and MC questions. Specifically, on MC questions, fine-tuning on axiomatic instances of d-separation yields a 17% improvement over the base model, even surpassing GPT-4. These strong zero-shot improvements across different question types and graph structures—notably outperforming larger models like GPT-4, indicate the effectiveness of axiomatic fine-tuning in enhancing causal reasoning.
@@ 7.2 Evaluation on Corr2Cause Dataset
Llama-3-8b-Instruct 0.11 0.15 0.08 0.77
Llama-3-8b-Instruct 0.23 0.16 0.46 0.54
fine-tuned Dsep (175k)
Llama-3-8b-Instruct 0.32 0.20 0.93 0.39
fine-tuned Transitivity (175k)
GPT-4 (from (Jin et al., 2024)) 0.29 0.21 0.48 0.64
Table 4: Evaluation on the Corr2Cause Task from (Jin et al., 2024). We fine-tune Llama-3-8B on our transitivity and d-separation datasets and evaluate on the Corr2Cause dataset in a zero-shot setting. We observe a significant increase in performance (over 20 p.p. improvement in F1 score compared to the base Llama model).
CLEAR D-Separation task (YN)
Models Accuracy
Llama-3-8B-Instruct 60.0
Llama-3-8B-Instruct fine-tuned 70.0
GPT-4 63.33
CLEAR D-Separation task (MC)
Models Accuracy
Llama-3-8B-Instruct 33.0
Llama-3-8B-Instruct fine-tuned 50.0
GPT-4 36.67
Table 5: Evaluation on CLEAR dataset (Chen et al., 2024) We fine-tune the LMs on our d-separation dataset and evaluate on the CLEAR dataset. We observe a significant increase in performance compared to the baseline, which highlights the efficacy of axiomatic training.
Dataset details. Corr2Cause (Jin et al., 2024) provides a more complex dataset to evaluate models on different causal tasks. Each data instance in the benchmark includes correlational relationships described in natural language for graphs with 3 to 6 nodes; the goal is to infer the truth value of a hypothesis. The hypothesis consists of six different kinds of graphical relationships between pairs of variables: Parent, Ancestor, Child, Descendant, Collider, and Confounder. This task is significantly harder than applying a single axiom. First, one needs to infer the causal graph from the correlation statements. This requires knowledge of d-separation and Markov condition (Appendix E.1). Then, one needs to use the transitivity axiom to identify the direct and indirect relationships to identify the children, ancestors, colliders, and confounders in the causal graph.
It was found that popular LLMs fail to perform well on this benchmark. While fine-tuning on this dataset shows promise, the fine-tuned language models also fail in out-of-distribution settings generated by simply perturbing these questions or changing node names. As for the CLEAR benchmark, we will test our axiomatically fine-tuned models exactly in the out-of-distribution setting: the models have only been trained on synthetic axiomatic data (described in Section 4.1) and not seen any instance of Corr2Cause. Specifically, we fine-tune two models, one based on transitivity data and one based on d-separation data, given that the questions involves multiple axioms. Due to the high imbalance between labels in the dataset, we use F1 score to measure performance.
Results. Our results highlight Llama-3-8b-Instruct Base model’s poor performance on this task, while axiomatic fine-tuning results in significant performance gains of around 20 p.p. (Table 4). Fine-Tuning on both transitivity and d-separation improves performance. Notably, transitivity fine-tuning led to the largest gains, even surpassing GPT-4, likely due to its direct relevance in inferring graph relationships required for the hypothesis covered in the Corr2Cause benchmark (such as identifying child, ancestor, parent, descendant). The significant gains in F1 scores under an out-of-distribution setting demonstrate the effectiveness of axiomatic fine-tuning.
@@ 8 Discussion and Conclusion
We introduced a general framework, axiomatic training, to add axioms and simple rules of causality as inductive priors in ML models and showed its effectiveness in improving performance on downstream causal reasoning tasks. For applicability to real-world tasks, we also showed that it can be used to finetune pre-trained language models. Our study shows the importance of diversity in axiomatic data to enhance generalizability.
Generalization to Logical Reasoning. While our axiomatic training approach focuses on causal reasoning, it can be applied to any formal system such as deductive logical reasoning. Recent work (Saparov et al., 2023) highlights LLMs’ deterioration in performance as reasoning depth increases. Given the similarity between our setup and such tasks, it would be interesting to explore if axiomatic training can improve deductive reasoning in language models. For example, transformers may be trained on logical axioms to improve language model’s reasoning; Morishita et al. (2024) provide promising evidence in this direction.
Verifiers for Causal Reasoning. Another application of axiomatic training may be in building verifiers that operate over language models’ output. For instance, our axiomatically finetuned models can be used to detect violations of causal rules such as transitivity and d-separation in a given text, and give inference-time feedback to the language model to improve its generation.
Implications for Training Language Models. Incorporating causal axiom demonstrations into pretraining or finetuning can boost reasoning in small language models, enabling models like Phi-3 to reach GPT-4-level accuracy on causal tasks. Papadimitriou & Jurafsky (2023) show that pretraining on synthetic formal languages introduces beneficial inductive biases. Similarly, pretraining on synthetic axiomatic data may enhance language models’ reasoning abilities. Another promising direction is to use axiomatic training to enhance base language models before applying post-training with reinforcement learning (RL). Since RL-based alignment typically amplifies the capabilities of a base model, incorporating causal understanding through axiomatic training could provide a stronger foundation (Gandhi et al., 2025).
@@
This paper presents work whose goal is to advance the field of Machine Learning, by proposing effective training strategies to improve causal reasoning capabilities of language models. Robust causal reasoning has clear upside: decision-support tools in medicine, public policy, and the social sciences may become more trustworthy under distribution shifts and easier for experts to scrutinize. We also recognize possible downsides, flawed or adversarially fine-tuned models could generate spurious causal claims that reinforce bias or justify harmful interventions. To mitigate these risks, we release code and synthetic training data openly, detail known limitations, and encourage independent replication and auditing. On balance, we believe that equipping language models with transparent, principled causal reasoning capabilities will yield net benefits for society while enabling informed oversight.
@@ A.1 Additional Results: Role of Data Diversity and Positional Encoding
Appendix B Zero Shot Setting
To evaluate the baseline models, we follow a simple zero-shot prompting strategy. For each tuple, we provide the natural language expression of the causal graph (Premise) followed by the question (Hypothesis) and prompt the LM to answer it in either ‘Yes’ or ‘No’ (Label). Here is an example prompt: “EX causes T. T causes 9. 9 causes W. W causes 7. 7 causes M. M causes a. Does EX cause T? Answer in ‘Yes’ or ‘No’ only”. See Table LABEL:tab:prompt_table contains examples of prompts used.
Appendix C Multi-Shot Prompt
C.1 Cause-Effect Inference Task
@@
Appendix D Results and Analysis
@@
Table A5: Results on node name length generalization. TS1 and TS2 denote Training Data setup 1 and 2 from Sec. 4. OCC is the third data setup comprising of sequential causal chains. SPE: Sinusoidal PE, LPE: Learnable PE, w/o PE: No PE, RoPE: Rotary Position Embedding. Model remains the same across all setups (67 Million parameter based). For longer node names, NoPE performs best on sequential linear setup. Accuracy metric is used.
@@
Appendix E Preliminaries and Notations
@@
Let $\mathcal{M}=(\bm{X},\bm{U},\mathcal{F})$ be a causal model defined over a set of endogenous variables $\bm{X}$, exogenous variables $\bm{U}$ and the causal relationship between them is defined by a set of structural equations $\mathcal{F}$ (Galles & Pearl, 1997). Let $\mathcal{G}$ be the causal graph associated with the causal model $\mathcal{M}$ where the nodes $\bm{V}$ in $\mathcal{G}$ correspond to the variables in $\mathcal{M}$ and an edge $V_{i}\rightarrow V_{j}$ between any two nodes $V_{i},V_{j}$ denote the causal relationship between them. The causal relationship of node $X_{i}$ is characterized by the functional relationship $f_{i}\in\mathcal{F}$ s.t., $x_{i}=f_{i}(\bm{pa}_{i},\bm{u}_{i})$. Here $\bm{pa}_{i}$ are the parent of the node $X_{i}$ is the corresponding causal graph $\mathcal{G}$ and $\bm{u_{i}}\subseteq\bm{U}$ are set of exogenous variables influencing the endogenous variable $X_{i}$. In our work, we assume that there are no hidden confounders so we have one exogenous variable corresponding to every endogenous variable i.e. $\bm{u}_{i}=U_{i}$. Each exogenous variable has an associated probability distribution which quantifies the uncertainty in the system i.e. $u_{i}\sim\mathbb{P}(u_{i})$. Thus the joint distribution of the exogenous variable is given by $\mathbb{P}(\bm{U})$. Since any endogenous variable is a deterministic function of other endogenous and exogenous variables the probability distribution corresponding to the endogenous variable is the push-forward of the exogenous variable i.e $\mathbb{P}(\bm{X})\triangleq\mathbb{P}^{\#}(\bm{U})$.
E.1 Definitions
Following the formal definitions provided by (Jin et al., 2024), we explain the following terminologies: